OpenStudy (deadshot):

What is 42° 15' 5" in in decimal degrees?

OpenStudy (anonymous):

this one isn't too bad if you understand this idea: ' means minutes, so 60 minutes = 1 degree and "" means seconds, so 60 seconds = 1 minute so start with the last number first, 5" so 5/60 = 0.08333 which is now the same as saying 42degrees 15.083333 minutes make sense?

OpenStudy (anonymous):

\[42+\frac{15}{60}+\frac{5}{3600}=\frac{30421}{720}=42.2514 \]

OpenStudy (deadshot):

Thanks!